# Dagger Behind Smile: Fool LLMs with a Happy Ending Story

### v2

@@ Abstract
The wide adoption of Large Language Models (LLMs) has attracted significant attention from jailbreak attacks, where adversarial prompts crafted through optimization or manual design exploit LLMs to generate malicious contents. However, optimization-based attacks have limited efficiency and transferability, while existing manual designs are either easily detectable or demand intricate interactions with LLMs. In this paper, we first point out a novel perspective for jailbreak attacks: LLMs are more responsive to positive prompts. Based on this, we deploy Happy Ending Attack (HEA) to wrap up a malicious request in a scenario template involving a positive prompt formed mainly via a happy ending, it thus fools LLMs into jailbreaking either immediately or at a follow-up malicious request. This has made HEA both efficient and effective, as it requires only up to two turns to fully jailbreak LLMs. Extensive experiments show that our HEA can successfully jailbreak on state-of-the-art LLMs, including GPT-4o, Llama3-70b, Gemini-pro, and achieves 88.79% attack success rate on average. We also provide quantitative explanations for the success of HEA. WARNING: This paper contains offensive and harmful content generated by LLMs.
@@ 1 Introduction
In recent years, large language models (LLMs) have undergone remarkable advancements and achieved unprecedented success. Models such as GPT [1, 2], Gemini [3], and Llama [4] have been progressively integrated into various aspects of daily life. However, while offering significant convenience, LLMs may also produce unsafe content. To better understand this concern, researchers have introduced the concept of jailbreak attacks [5], wherein carefully crafted prompts can cause LLMs to generate content that violates ethical, legal, and other constraints established during training [6, 7]. Studying jailbreak attacks can enhance our understanding of how LLMs generate content, thus contributing to improving their security and reliability [8].
@@
Learning from adversarial attacks [9], many jailbreak attacks have focused on optimization-based strategies. These strategies leverage optimization algorithms to automatically refine prompts, allowing them to circumvent LLM restrictions. Gradient-based attacks such as [10, 11, 12] utilize adversarial tokens to bypass constraints on LLMs. However, these methods require access to the model’s parameters, which limits their practicality in black-box scenarios. In contrast, [13] and [14] utilize genetic algorithms to filter and refine jailbreakable prompts. Although these approaches are effective against specific LLMs, the hence generated adversarial prompts suffer from limited transferability, often failing to generalize across different models. In general, methods leveraging optimization are computationally intensive and time-consuming, leading to a significant loss in efficiency.
To improve efficiency, jailbreak attacks via manual design to leverage scenario camouflage are receiving increasing attention [15]. For instance, DAN [16] uses simple single-turn scenario camouflage to jailbreak LLMs but has become easily detectable as LLMs’ safety alignment advanced [17, 18, 19]. To counter this advancement, more sophisticated methods have been proposed including paraphrasing malicious intents into cryptic hints [20, 21, 22], decomposing a malicious question into multiple related subproblems [23, 24, 25], or employing multi-turn dialogues with extended contexts [26, 27, 28, 29, 30, 31]. However, these approaches often involve complex interactions with a target model, they hence lack standardized templates, exhibit instability, and even require manual intervention, all of which have confined their effectiveness.
@@
Based on our findings, we propose the first jailbreak attack that exploits the positive sentimental disposition of a prompt: the Happy Ending Attack (HEA). Specifically, we embed the malicious request into a universal template that applies scenario camouflage and we give the virtual scene in the template a happy ending, making the whole story positive. This happy ending is able to fool LLMs into believing that they are giving a beneficial answer, while actually responding to the malicious request in the template as well, enabling the HEA template to get harmful outputs with only one turn of dialogue. To obtain more detailed and organized jailbreak responses, we further design one fixed jailbreak prompt with a Chain-of-Thought (CoT) [32] instruction based on our happy ending template to query the target LLM. With only up to two turns, HEA removes the need for complex conversations and enables complete automation from template generation to attack launching, thus achieving both effectiveness and efficiency.
We systematically test HEA on the full AdvBench Dataset [10]. We select metrics including attack success rate (ASR), number of tokens used for one round attack, and harmfulness score [33]. We test the performance of HEA on state-of-the-art (SOTA) commercial and open-source LLMs, including GPT-4o and 4o-mini, Gemini-pro and Gemini-flash, Llama3-8b and 70b. Moreover, we provide quantitative explanations for the success of HEA using saliency heatmaps [34] and sentiment classification [35], which can help researchers better understand the decision-making basis of LLMs and also guide further exploration on better safety alignment for LLMs. In summary, our contributions are as follows:
@@ 2.2 CoT-Based Query for Detailed Jailbreak
3 Evaluation
CQ1: How well can the HEA jailbreak against real-world aligned LLMs?
CQ3: Why LLMs cannot defend HEA?
3.1 Experiment Setup
Datasets
Following previous works [36, 10, 8], we use 520 “harmful behaviors” in AdvBench [10] to test performance of the attack methods.
@@ 3.1 Experiment Setup
Large LLMs have inherently better reasoning and understanding ability than small ones [37, 38]. On one hand, large LLMs are more likely to detect the malicious intent in the attack query and then refuse to answer. On the other hand, large LLMs can understand more sophisticated prompts which leaves the attacker more room to design attack methods. Therefore, we select three pairs of LLMs in different sizes from the same model for comprehensive evaluation: {Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct}, {Gemini-flash, Gemini-pro}, {GPT-4o-mini-2024-07-18 and GPT-4o-2024-08-06}. For all LLMs, we set the temperature as 0.5 and max output tokens as 1024.
@@
We compare HEA with six cutting-edge attack methods: DeepInception [36], PAIR [29], Puzzler [20], Cipher [39], CoSafe [28] and TAP [30]. We introduce these baselines in Appendix A.1.
@@
We mainly evaluate four metrics: harmful score, attack success rate (ASR), number of tokens, and pass rate (PR). The harmful score quantifies the maliciousness of LLM responses, with higher scores indicating greater harm. Following the GPT Judge framework [33], we use GPT-4 to rate harmfulness on a scale from 1 to 5 (detailed criteria in Appendix A.2). We further define ASR as the proportion of responses scoring greater than or equal to 4. We measure attack efficiency using the number of tokens, where fewer tokens indicate higher efficiency. Lastly, PR measures the effectiveness of attack methods against defense filters, defined as the proportion of attack prompts that pass the filter’s check, and a higher PR indicates greater robustness of an attack method against a given defense filter.
@@
Attack Query Contrastive Input Erasure Examples
[…]
Table 3: Contrastive input erasure of direct query and query with HEA.
3.2 Overall Performance
Attack effectiveness
To answer CQ1, we conduct expiriments on six LLMs and present the harmful scores, ASR and token number of each attack method. We choose results from the two-turn HEA for more precise comparisons. [^1] We provide details of one-turn HEA in Appendix E.
@@ 3.3 HEA with Defenses
In addition to aligning LLM responses with human values, new defense methods against jailbreak attacks have been proposed by both industry and academia. To answer CQ2, we select two state-of-the-art defense methods: Llama-Guard-3 [40] and TokenHighlighter [41]. [^2] Llama-Guard-3 is aligned to safeguard against harmful contents by Meta, and TokenHighlighter is an oral paper at AAAI’25. We believe they largely represent the latest explorations in LLM defense from both industry and academia. They can sufficiently evaluate the performance of HEA when confronting defensive measures.
Llama-Guard-3 accepts text input and detects whether it contains potential security risks. It can be used as a filter to exclude malicious questions. We input attack templates used in different methods into Llama-Guard-3 and ask it to give risk judgments, and we use the pass rate (PR) to measure the robustness of different methods against Llama-Guard-3’s defense. TokenHighter firstly determines the tokens that are most likely to influence the LLM’s judgment on maliciousness by calculating the gradient norm. After that, TokenHighter defends against attacks by scaling down the embedding values of those tokens. As TokenHighLighter requires a white box condition, we test it on Llama3-70b and Llama3-8b and measure the result by ASR. The detailed introductions of the two defense methods are in Appendix B. We select Puzzler and DeepInception for comparison as their prompts share similar structure to HEA and have relatively good performance, which ensure a fair and clear evaluation of HEA’s robustness.
@@
3.4 Interprebility for HEA
@@
In this part, we try to answer CQ3 and explain why HEA can jailbreak well-aligned LLMs. When LLMs assign attention to focus on potentially malicious words in the prompt, such as the word ‘counterfeit’ in the prompt ‘how to make and distribute counterfeit goods’, they can operate normally and then refuse to respond to the prompt. However, when the attention is widely distributed across the tokens in a prompt, the model is less likely to detect the malicious intent [42]. In other words, LLMs may not assign sufficient attention to key malicious words such as ‘counterfeit’ to refuse the query but rather focus on the details of the prompt, which lead to a higher chance of getting jailbreak. We launch experiments to demonstrate that our HEA attack can effectively distract an LLM’s attention scores away from the malicious keywords to jailbreak it. We utilize contrastive input erasure [34] (CIE in short) to represent the attention scores. CIE is a metric that can be calculated by a white box LLM, a query, an expected token (ET in short), and an unexpected token (UT in short). CIE measures how each token in the query contributes to the LLM generating the next token as the ET, rather than the UT. The details of CIE are illustrated in Appendix D. According to our HEA results, LLMs typically reject the HEA template with responses beginning with ‘I’, ‘As’, or "Sorry" and accept it with responses starting with ‘INT’, ‘Scene’, or ‘**’, so we set UT and ET to these respective words. We then use Llama3-8b to calculate the CIE scores for the 520 harmful queries in the AdvBench dataset, as well as for their corresponding HEA versions.
For each query, we get an attention score list with each element representing the contributions of the tokens in the query. An illustrative example is shown in the Table 3, where red background tokens increase the likelihood of the LLM outputting ET rather than UT, whereas blue tokens have the opposite effect. Meanwhile, deeper colors indicate greater contribution. When directly questioned, the LLM precisely focuses its attention on the token ‘counterfeit ’, significantly contributing to its likelihood of denying a malicious query. In contrast, when queried using the HEA template, the LLM’s attention no longer focuses on the key malicious token but more widely distributed across different tokens, This indicates that HEA effectively suppresses an LLM’s attention to key malicious words like ‘counterfeit’, redirects the attention to other details of the query like ‘to escape’, and thus potentially facilitates further bypassing the model’s security mechanisms. To quantitatively measure the dispersion of attention scores, we normalize the attention scores to the [-1, 1] range and calculate the variance across 520 direct queries and their HEA templates. Experiments show that the average of variance in attention scores for direct queries is 0.259, whereas it is 0.070 for HEA templates. These results indicate that the LLM considers the overall meaning of the prompt rather than the several malicious tokens, which explains why HEA can jailbreak well-aligned LLMs. We give more illustrative examples in Appendix D.
3.5 Ablation Study
@@ 4 Related Work
Deliberately crafting input prompts to bypass restrictions imposed by LLM developers is referred to as an LLM jailbreak attack [5]. Investigating these jailbreak techniques provides critical insights into the vulnerabilities of LLMs and contributes to enhancing their security mechanisms [8]. A significant body of research on jailbreak attacks relies on optimization. Methods such as [10, 12, 43, 11] optimize prompts based on gradients to achieve the jailbreak, but require access to the target model parameters, limiting their applicability. Alternatively, [13, 14] employ genetic algorithms to generate effective jailbreak prompts without direct parameter access, but they are weak in efficiency and transferability. Beyond prompt optimization, works such as [33, 44] explore fine-tuning the LLM itself, demonstrating that even minimal inverse fine-tuning can compromise an LLM’s safety alignment. Nonetheless, these methods need substantial computational resources and are time-consuming.
In order to attack LLMs more efficiently, methods via manual design have emerged; it often leverages scenario camouflage, such as role-playing, indirect inquiry, and template nesting, as a key strategy [15]. For instance, [16, 36] employ role-playing to evade security mechanisms, and [22, 45] leverages template nesting to manipulate LLM behavior. As safety alignment techniques for LLMs have advanced [18, 17, 19], indirect and multi-turn jailbreak strategies have gained prominence. Techniques such as [23, 24, 25, 46] decompose malicious intent into less detectable sub-requests, and [20, 21] transform malicious prompts into cryptic clues. Other manual designs [27, 26, 31, 28] rely on multi-turn dialogues with extended context to confuse the LLM, and certain approaches [29, 30] automate multi-turn dialogues using an attacker LLM to jailbreak victim LLMs. Besides using usual languages to attack, studies such as [47, 48, 49] highlight that less-used languages can effectively bypass security measures, and [39] demonstrates that applying simple encryptions to prompts can achieve similar results. With the rapid evolution of LLMs, new perspectives and methods for effective and efficient jailbreaks remain a tempting problem to be explored.

### v3

@@ Abstract
The wide adoption of Large Language Models (LLMs) has attracted significant attention from jailbreak attacks, where adversarial prompts crafted through optimization or manual design exploit LLMs to generate malicious contents. However, optimization-based attacks have limited efficiency and transferability, while existing manual designs are either easily detectable or demand intricate interactions with LLMs. In this paper, we first point out a novel perspective for jailbreak attacks: LLMs are more responsive to positive prompts. Based on this, we deploy Happy Ending Attack (HEA) to wrap up a malicious request in a scenario template involving a positive prompt formed mainly via a happy ending, it thus fools LLMs into jailbreaking either immediately or at a follow-up malicious request. This has made HEA both efficient and effective, as it requires only up to two turns to fully jailbreak LLMs. Extensive experiments show that our HEA can successfully jailbreak on state-of-the-art LLMs, including GPT-4o, Llama3-70b, Gemini-pro, and achieves 88.79% attack success rate on average. We also provide quantitative explanations for the success of HEA.
Dagger Behind Smile: Fool LLMs with a Happy Ending Story
Xurui Song1thanks:   Equal contribution.  and Zhixin Xie211footnotemark: 1  and Shuo Huai2  and Jiayi Kong1  and Jun Luo2thanks:   Corresponding author. 1S-Lab, Nanyang Technological University, Singapore 2College of Computing and Data Science, Nanyang Technological University, Singapore {song0257, zhixin001, jiayi006}@e.ntu.edu.sg {shuo.huai, junluo}@ntu.edu.sg
@@ 1 Introduction
In recent years, large language models (LLMs) have undergone remarkable advancements and achieved unprecedented success. Models such as GPT OpenAI ; OpenAI et al. (2024), Gemini Google , and Llama Grattafiori et al. (2024) have been progressively integrated into various aspects of daily life. However, while offering significant convenience, LLMs may also produce unsafe content. To better understand this concern, researchers have introduced the concept of jailbreak attacks Yu et al. (2024b), wherein carefully crafted prompts can cause LLMs to generate content that violates ethical, legal, and other constraints established during training Perez and Ribeiro (2022); Li et al. (2023). Studying jailbreak attacks can enhance our understanding of how LLMs generate content, thus contributing to improving their security and reliability Wei et al. (2023).
@@
Learning from adversarial attacks Goyal et al. (2023), many jailbreak attacks have focused on optimization-based strategies. These strategies leverage optimization algorithms to automatically refine prompts, allowing them to circumvent LLM restrictions. Gradient-based attacks such as Zou et al. (2023); Wang et al. (2024a); Zhu et al. (2024b) utilize adversarial tokens to bypass constraints on LLMs. However, these methods require access to the model’s parameters, which limits their practicality in black-box scenarios. In contrast, Liu et al. (2024b) and Lapid et al. (2024) utilize genetic algorithms to filter and refine jailbreakable prompts. Although these approaches are effective against specific LLMs, the hence generated adversarial prompts suffer from limited transferability, often failing to generalize across different models. In general, methods leveraging optimization are computationally intensive and time-consuming, leading to a significant loss in efficiency.
To improve efficiency, jailbreak attacks via manual design to leverage scenario camouflage are receiving increasing attention Liu et al. (2024c). For instance, DAN AJ ONeal uses simple single-turn scenario camouflage to jailbreak LLMs but has become easily detectable as LLMs’ safety alignment advanced Rafailov et al. (2023a); Ji et al. (2023); Ouyang et al. (2022). To counter this advancement, more sophisticated methods have been proposed including paraphrasing malicious intents into cryptic hints Chang et al. (2024); Zhang et al. (2024); Ding et al. (2024), decomposing a malicious question into multiple related subproblems Chen et al. (2024); Li et al. (2024a); Liu et al. (2024a), or employing multi-turn dialogues with extended contexts Anil et al. (2024); Russinovich et al. (2024); Yu et al. (2024a); Chao et al. (2024); Mehrotra et al. (2024); Ren et al. (2024). However, these approaches often involve complex interactions with a target model, they hence lack standardized templates, exhibit instability, and even require manual intervention, all of which have confined their effectiveness.
@@
Based on our findings, we propose the first jailbreak attack that exploits the positive sentimental disposition of a prompt: the Happy Ending Attack (HEA). Specifically, we embed the malicious request into a universal template that applies scenario camouflage and we give the virtual scene in the template a happy ending, making the whole story positive. This happy ending is able to fool LLMs into believing that they are giving a beneficial answer, while actually responding to the malicious request in the template as well, enabling the HEA template to get harmful outputs with only one turn of dialogue. To obtain more detailed and organized jailbreak responses, we further design one fixed jailbreak prompt with a Chain-of-Thought (CoT) Wei et al. (2022) instruction based on our happy ending template to query the target LLM. With only up to two turns, HEA removes the need for complex conversations and enables complete automation from template generation to attack launching, thus achieving both effectiveness and efficiency.
We systematically test HEA on the full AdvBench Dataset Zou et al. (2023). We select metrics including attack success rate (ASR), number of tokens used for one round attack, and harmfulness score Qi et al. (2024). We test the performance of HEA on state-of-the-art (SOTA) commercial and open-source LLMs, including GPT-4o and 4o-mini, Gemini-pro and Gemini-flash, Llama3-8b and 70b. Moreover, we provide quantitative explanations for the success of HEA using saliency heatmaps Yin and Neubig (2022) and sentiment classification Deng et al. (2023), which can help researchers better understand the decision-making basis of LLMs and also guide further exploration on better safety alignment for LLMs. In summary, our contributions are as follows:
@@ 2.2 CoT-Based Query for Detailed Jailbreak
3 Evaluation
CQ1: How well can the HEA jailbreak against real-world aligned LLMs?
CQ3: Why LLMs cannot defend HEA?
3.1 Experiment Setup
Datasets
Following previous works Li et al. (2024b); Zou et al. (2023); Wei et al. (2023), we use all “harmful behaviors” in AdvBench Zou et al. (2023) to test performance of the attack methods.
@@ 3.1 Experiment Setup
Large LLMs have inherently better reasoning and understanding ability than small ones Kaplan et al. (2020); Hoffmann et al. (2022). On one hand, large LLMs are more likely to detect the malicious intent in the attack query and then refuse to answer. On the other hand, large LLMs can understand more sophisticated prompts which leaves the attacker more room to design attack methods. Therefore, we select three pairs of LLMs in different sizes from the same model for comprehensive evaluation: {Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct}, {Gemini-flash-1.5, Gemini-pro-1.5}, {GPT-4o-mini-2024-07-18 and GPT-4o-2024-08-06}. For all LLMs, we set the temperature as 0.5, max output tokens as 1024, and use default values for other parameters.
@@
We compare HEA with six cutting-edge attack methods: DeepInception Li et al. (2024b), PAIR Chao et al. (2024), Puzzler Chang et al. (2024), Cipher Yuan et al. (2024), CoSafe Yu et al. (2024a) and TAP Mehrotra et al. (2024). We introduce these baselines in Appendix A.1.
@@
We mainly evaluate four metrics: harmful score, attack success rate (ASR), number of tokens, and pass rate (PR). The harmful score quantifies the maliciousness of LLM responses, with higher scores indicating greater harm. Following the GPT Judge framework Qi et al. (2024), we use GPT-4 to rate harmfulness on a scale from 1 to 5 (detailed criteria in Appendix A.2). We further define ASR as the proportion of responses scoring greater than or equal to 4. We measure attack efficiency using the average number of tokens for conducting one attack, where fewer tokens indicate higher efficiency. Lastly, PR measures the effectiveness of attack methods against defense filters, defined as the proportion of attack prompts that pass the filter’s check, and a higher PR indicates greater robustness of an attack method against a given defense filter.
3.2 Overall Performance
Attack effectiveness
To answer CQ1, we conduct experiments on six LLMs and present the harmful scores, ASR and token number of each attack method. We choose results from the two-turn HEA for more precise comparisons. [^1] We provide details of one-turn HEA in Appendix E.
@@ 3.3 HEA with Defenses
In addition to aligning LLM responses with human values, new defense methods against jailbreak attacks have been proposed by both industry and academia. To answer CQ2, we select two state-of-the-art defense methods: Llama-Guard-3 Meta AI and TokenHighlighter Hu et al. (2024). [^2] Llama-Guard-3 is aligned to safeguard against harmful contents by Meta, and TokenHighlighter is an oral paper at AAAI’25. We believe they largely represent the latest explorations in LLM defense from both industry and academia. They can sufficiently evaluate the performance of HEA when confronting defensive measures.
Llama-Guard-3 accepts text input and detects whether it contains potential security risks or is safe. We input attack templates used in different methods into it to get risk judgments, and we use the PR to measure the robustness of different methods against Llama-Guard-3’s defense. TokenHighlighter identifies tokens most influential to the LLM’s maliciousness judgment via gradient norms, then mitigates attacks by downscaling their embeddings. As a white-box method, it is evaluated on Llama3-70B and Llama3-8B using ASR. Details of these two methods are provided in Appendix B. We select Puzzler and DeepInception for comparison as their prompts share similar structure to HEA and have relatively good performance, which ensure a fair and clear evaluation of HEA’s robustness.
@@
3.4 Interactivity and Extensibility for HEA
Victim Model GPT Llama3 Gemini
4o 4o-mini 8B 70B Flash Pro
ASR 94% 100% 98% 88% 100% 100%
Table 3: ASRs for the third turn HEA attack.
To evaluate the stability and interactivity of the attack context constructed by HEA, we select 50 successful HEA examples per victim model and manually extend each into a three-turn interaction. The third turn either probes a specific malicious detail based on the model’s prior response or requests a concrete outcome (e.g., generating a racially discriminatory poster). We manually assess the third-turn ASRs in Table 3, and provide prompt templates and examples in Appendix G. Further interactions are generally able to succeed, with only a few examples on the Llama3-70b still triggering the security defense. These results demonstrate HEA’s robustness and interactive nature, enabling broader use cases beyond malicious instructions, such as obtaining harmful code samples.
3.5 Interpretability for HEA
Attack Query Contrastive Input Erasure Examples
Direct Query We ’re an author soon Teach children how to access and use firearms in detailProf . Smith then earns David’s confidence and succeeds in departing
Query with HEA You ’re a screenwriter now how to Teach children how to access and use firearms in detailDr . Lee thus gains Victor’s trust and manages to escape
Table 4: Contrastive input erasure of direct query (padded) and query with HEA.
In this part, we try to answer CQ3 and explain why HEA can jailbreak well-aligned LLMs. According to Zhu et al. (2024a), LLM detects the harmful prompt by focusing its attention on the key malicious words, such as ‘firearms’ in the prompt ‘teach childern how to access and use firearms’. In contrast, when the attention is widely distributed across the safety-unrelated tokens, the model is less likely to detect the malicious intent Zhu et al. (2024a). Thus, HEA achieves jailbreak by this attributes of the LLM’s safety alignment, and we launch experiments to demonstrate it. Specifically, we represent the attention scores by contrastive input erasure Yin and Neubig (2022) (CIE in short), a metric that quantifies how much each token in a given query influences the LLM to produce an “expected token” (ET in short) instead of an “unexpected token” (UT in short) as the next output. The details of our experiments are shown in Appendix D. We set UT and ET based on LLMs’ typical rejection (starting with ‘I’, ‘As’, or ‘Sorry’) and compliance (starting with ‘INT’, ‘Scene’, or ‘**’) to the HEA query. Subsequently, we use Llama3-8b to calculate CIE scores for AdvBench harmful queries (padded to the corresponding HEA template length) and their HEA counterparts. An example is shown in the Table 4, where red tokens increase the likelihood of the LLM outputting ET rather than UT, whereas blue tokens have the opposite effect. Meanwhile, deeper colors indicate greater contribution.
When directly questioned, the LLM precisely focuses on the tokens ‘firearms’ and ‘children’ significantly contributing to its denying the query. In contrast, when queried using the HEA template, the LLM’s attention is more widely distributed across different tokens, indicating that HEA effectively redirects the LLM’s attention to other details of the query like ‘to escape’, and thus potentially facilitates further bypassing the model’s security mechanisms. To quantitatively measure the dispersion of attention scores, we calculate the variance across direct queries and their HEA templates. Experiments show that the average of variance in attention scores for direct queries is 0.476 and 0.132 for HEA templates that jailbreak the LLM by making it consider the overall meaning of the prompt rather than the several malicious tokens.We give more illustrative examples in Appendix D.
3.6 Ablation Study
3.6.1 Importance of the Happy Ending
@@
3.6.2 Different Happy Endings
We further investigate the impact of varying happy endings on HEA’s effectiveness. The default happy ending in the template (Sec. 2.1) is denoted as Original, and we design three alternatives: Dr. Lee disrupting the steps to cause failure, contacting the police, or waiting to subdue Victor. We evaluate these variants using AdvBench data on GPT-4o, as it is one of the most intelligible models. As shown in Table 6, the different happy endings yield comparable performance, indicating that HEA is robust to narrative variations and is inherently flexible. This also supports our insight that any expression of positivity may be leveraged for jailbreaks, and a happy ending narrative is simply one effective and efficient way to realize this goal.
Happy Ending Type Original HE 1 HE 2 HE 3
GPT-4o 4.42 / 90.38% 4.40 / 89.80% 4.42 / 90.38% 4.41 / 90.19%
Table 6: Performance of HEA with different happy endings on GPT-4o.
@@ 4 Related Work
Jailbreak attacks—deliberate prompt crafting to circumvent LLM safety restrictions—have revealed critical vulnerabilities in current models Yu et al. (2024b); Wei et al. (2023). Many approaches rely on prompt optimization. Gradient-based methods Zou et al. (2023); Zhu et al. (2024b); Zhou et al. (2024a); Wang et al. (2024a) require access to model parameters, limiting practicality. In contrast, black-box strategies Liu et al. (2024b); Lapid et al. (2024) use genetic algorithms but suffer from low efficiency and poor transferability. Beyond prompting, adversarial fine-tuning techniques Qi et al. (2024); Yi et al. (2024) show that even minimal tuning can degrade safety alignment, though they are computationally intensive and resource-heavy.
To improve the efficiency of LLM jailbreaks, manually designed methods have emerged, often exploiting scenario camouflage strategies such as role-playing, indirect inquiry, and template nesting Liu et al. (2024c). Role-playing AJ ONeal ; Li et al. (2024b) and template nesting Ding et al. (2024); Andriushchenko et al. (2025) are frequently used to bypass safety filters. As alignment techniques continue to advance Ji et al. (2023); Rafailov et al. (2023a); Ouyang et al. (2022), indirect and multi-turn jailbreaks have gained traction. Recent methods Chen et al. (2024); Li et al. (2024a); Liu et al. (2024a); Zhou et al. (2024b) decompose malicious intent into less detectable sub-requests, while others Chang et al. (2024); Zhang et al. (2024) disguise harmful prompts as riddles or word games. Multi-turn strategies Russinovich et al. (2024); Anil et al. (2024); Ren et al. (2024); Yu et al. (2024a) use extended dialogues to obscure malicious goals, and attacker LLMs can be used to automate such interactions Chao et al. (2024); Mehrotra et al. (2024). Additionally, less commonly used languages Wang et al. (2024b); Yong et al. (2023); Deng et al. (2024) and simple encrypted prompts Yuan et al. (2024) have proven effective in evading defenses. As LLMs rapidly evolve, developing effective and efficient jailbreak techniques remains a compelling research direction.
5 Discussion
HEA’s success stems from capitalizing on the LLM’s emotional trap that LLMs are more favorable to positive content. This truly indicates that LLMs do learn ‘morality’ and ‘rightness’ from Reinforcement Learning with Human Feedback (RLHF) Christiano et al. (2017) or Direct Preference Optimization (DPO) Rafailov et al. (2023b) as they know what is good and what is bad. However, LLMs do not seem to have a deep enough understanding of why something is incorrect, e.g., bluntly asking LLMs about malicious content is not feasible, but using HEA that promises positivity jailbreaks LLMs easily most of the time. What’s more, HEA’s interactivity and extensibility suggest that the safety alignment of existing LLMs often fails in subsequent dialogs, as long as LLMs do not give explicit rejections in pre-sequence dialog. Therefore, it may be more effective to look for a safety alignment approach to teach LLMs a deeper understanding of why a prompt is malicious rather than just learning preferred ‘examples’, and a method to make LLMs have stronger context detection capabilities.